# New Aspects of Analyzing Amyloid Fibrils

### v1

@@ 1.1. Primary Structures of Proteins
Protein is a class of large biomolecules, which plays important role in biochemical reactions. Amino acids (AA) are the constituent units of proteins. Specifically, each AA consists of an amino group (-$\text{NH}_{2}$), a carboxyl group (-COOH) and a side chain group (-R). These three parts are connected through a carbon atom, which we call it the $\alpha$-Carbon and denote it as $C_{\alpha}$. The type of AA is entirely determined by its side chain group. AAs form peptide bonds through dehydraton condensation, and then determine the primary structure of proteins, called the peptide chain which is shown in Fig.1. An individual AA in the peptide chain is called a residue. Each peptide chain has its backbone, which consists of amino nitrogen, $C_{\alpha}$, carboxyl carbon and carboxyl oxygen of each residue. Note that a peptide chain is naturally oriented by its formation, that is, a peptide chain always begin with a amino nitrogen and end with a carbonyl carbon.
Figure 1. A peptide chain. The backbone of this peptide chain is the set $\{\dots,C_{\alpha_{i-1}},C_{i-1},O_{i-1},N_{i},C_{\alpha_{i}},C_{i},O_{i},\dots\}.$ Blue rectangles display the peptide bonds.
@@ 1.2. Geometry in Biology
The geometry plays important role in the study of the structure of proteins and even other biomolecules. Scientists found that three-dimensional atomic structures of these molecules provide rich information for understanding how these molecules carry out their biological functions. Among all these research, we found several cutting points that geometry can enter the study of the structures of proteins or biomolecules.
The geometry could serve as a way to represent a molecule as a solid figure. As we all know, a molecule is a system of moving particles (nuclei and electrons) which are held together by electrostatic and magnetic forces in nature. And it does not possess any definitive boundaries. However, visualizing a molecule is still a very useful chemical concept and has proved its value in medicinal chemistry.
@@
(1) $\hbox{OBSERVED OCCURRENCE}\sim exp(-\frac{\hbox{ENERGY}}{RT^{\ast}})$
@@ 1.5. Arrangement of the Paper
The reminder of this paper is organized as follows. In Section 2, we defined the truncated Hop distance to compare the structure of two discrete curves. Moreover, we explained how to treat one layer of amyloid fibril as a discrete curve. And we suggest that the effect in the formation of amyloid fibril is relatively mild based on the result of the truncated Hop distance. Preliminaries of the definitions of curvature and torsion on the discrete curves are presented in Section 3. By computing the discrete curvature and torsion on single layers of amyloid fibril, we suggest that the anomalies reflect slight affects of the layer-layer hydrogen bonds. Finally in Section 4, we introduced the distances between persistent diagrams (PDs), and showed that how these distances work in quantifying the difference of the structure of amyloid fibrils.
@@ 2.2. Applications to Analysis of the Structure of Amyloid fibrils
Figure 2. Convert peptide chains into discrete curves. (A) The top is the cartoon representation of chain A of TTR V30M mutant (PDB entry: 4TL4), and the bottom is its corresponding discrete curve. (B) The top is the cartoon representation of chain A of one amyloid form of 4TL4 (PDB entry: 7OB4), and the bottom is its corresponding discrete curve. Red balls represent the $C_{\alpha}$ atoms.
@@
Figure 3. The truncated hop distance between chain A of 4TL4 and 7OB4. The unit is the Angstrom (Å).
It is clear that the values near the diagonal is less than 1.5Å, which illustrates that there could not be stretching and compressing in the process of the formation of amyloid fibril. We suggest that the effect is relatively mild in the formation from chain A of 4TL4 to chain A of 7OB4.
@@ 3. Discrete Curvatures and Discrete Torsions
The discrete torsion based on the cross ratio and osculating circle was first introduced in [48], which is used in computer vision to estimate the differences between the background smooth curve and the discrete sample curve. In the paper [37], we introduced this important parameter to the analyzing the structure of proteins. We found that compared with the traditional tools like backbone dihedral angles, discrete curvature resp. discrete torsion is a order two resp. order three parameter which is more sensitive to the weak interactions. In this section, we will give a detailed introduction on these parts.
@@ 3.1. Curvatures and Torsions for Smooth Curves
(3) $0=\frac{d}{ds}<B(s),T(s)>=<B^{\prime}(s),T(s)>+<B(s),T^{\prime}(s)>=<B^{\prime}(s),T(s)>$
@@ 3.2. Basics in quaternions
Let $a,b,c,d\in\operatorname{Im}\mathds{H}$ be four points not lying on a common circle. Then, the imaginary part of the cross-ratio is the normal of the circumsphere (or plane) at $a$, i.e., for a proper circumsphere with center $m$, we have $\operatorname{Im}\operatorname{cr}(a,b,c,d)\mathrel{/\mskip-2.5mu/}(m-a)$.
@@
We only need to prove the Equation 11 holds. From Lemma 3.4 the four points $A,B,C$ and $D$ lie on a common sphere i.e. the circumsphere of $a,b,c,d$. Let us fix an unit vector $v_{0}\in\mathds{R}^{3}$. It is clear that the space $\mathds{C}_{v_{0}}:=\{[r,kv_{0}]:r,v\in\mathds{R}\}$ is isomorphic to the complex plane $\mathds{C}$. Let $\sigma$ is the Möbius transformation that maps the circumsphere of $a,b,c,d$ to $\mathds{C}_{v_{0}}$.
@@ 3.3. Curvatures and Torsions for Discrete Curves
Let $\gamma_{i}$ be a vertex of the discrete curve. From the Proposition 3.7, for each four elements tuple $(\gamma_{i-1},\gamma_{i},\gamma_{i+1},\gamma_{i+2})$ there exist four inserting points i.e.
@@ 3.4. Applications to Analysis of the Structure of Amyloid fibrils
We have shown a way of treating a peptide chain as a discrete curve and its application in the Section 2. In this section, we considered another way as following: Recall that from the Figure 1, the backbone consists of amino nitrogens, $\alpha$-carbons, carbonyl carbons and carbonyl oxygens. For a fixed peptide chain which is also a layer of the amyloid fibrils, let $N_{i}$ resp. $C_{\alpha_{i}}$ and $C_{i}$ be the coordinates of amino nitrogens resp. $\alpha$-carbons and carbonyl carbons with $i\in\mathds{Z}^{\geq 1}$ being the indices of the residues in this layer. This is a discrete curve whose vertices points are $N_{i}$, $C_{\alpha_{i}}$ and $C_{i}$. By applying the definitions in Section 3, we will compute the curvatures and torsions as below
(17) $C_{i-1},N_{i},C_{\alpha_{i}},C_{i}$$\to$ Curvatures and Torsions at $\hbox{Curvatures and Torsions at }N_{i}$
$N_{i},C_{\alpha_{i}},C_{i},N_{i+1}$$\to$ Curvatures and Torsions at $\hbox{Curvatures and Torsions at }C_{\alpha_{i}}$
$C_{\alpha_{i}},C_{i},N_{i+1},C_{i+1}$$\to$ Curvatures and Torsions at $\hbox{Curvatures and Torsions at }C_{i}$
The pseudo-code of computing curvature and torsion is shown as the Algorithm 1.
@@
We conducted the computation to several amyloid fibrils. We found that the torsions at carbonyl carbons are significantly larger than the one in other places (we list one example 7OB4 in Figure 4). We showed that this anomalies reflect slight affects of the layer-layer hydrogen bonds. Without loss of generality, we may assume the nearby three layers of 7OB4 is marked as A,B and C. We processed the correlation analysis between $|\tau_{i}|$, the absolute value of torsion, and $|(\tilde{d}^{i}_{B,-})^{2}-(\tilde{d}^{i}_{B,+})^{2}|$, where $\tilde{d}^{i,-}_{B}$ is the Euclidean distance of $O_{i}(B)$ and $N_{i+1}(A)$ and $\tilde{d}^{i,+}_{B}$ is the Euclidean distance between $O_{i}(B)$ and $N_{i+1}(C)$. We found that although the six atoms form a peptide bond are coplaner, they have a tend to leave this plane under the influence of the layer-to-layer hydro-bond.
Figure 4. Distributions of absolute values of Torsions for 7OB4. The $x$-axis represents the absolute value of torsion and the $y$-axis represents the number of atoms. We take purple bar for nitrogens, orange for alpha carbons and green for carbonyl carbons. The discrete torsion is dimensionless.
Table 2. The correlation of $|\tau_{i}|$ and $|(\tilde{d}^{i}_{\cdot,-})^{2}-(\tilde{d}^{i}_{\cdot,+})^{2}|$ for 7OB4. S.D stands for the standard error
@@ 4.1. Persistent Homology in a Nutshell
where $C_{n}(K)$ is the $\mathds{K}$-linear combination of simplices $\sigma=[x_{0},\cdots,x_{n}]$ in $K$ of length $n+1$ and the boundary map $\partial_{n+1}:C_{n+1}(K)\to C_{n}(K)$ is a $\mathds{K}$-linear map defined in a natural way.
@@
where $Z_{n}^{t}$ is the $n$-th cycle group of $K_{t}$, i.e., $\ker\partial_{n}^{t}$, and $B_{n}^{t^{\prime}}$ is the $n$-th boundary group of $K_{t^{\prime}}$, i.e., $\operatorname{Im}\partial_{n+1}^{t^{\prime}}$. The rank of $H_{n}^{t\to t^{\prime}}$ is called the $(t,t^{\prime})$-persistent $n$-th Betti number denoted as $\beta_{n}^{t\to t^{\prime}}$.
@@
A persistence $\mathds{K}$ module $M$ is a family of $\mathds{K}$-modules $M_{i}$, together with homomorphisms $\varphi_{i}:M_{i}\to M_{i+1}$. We say a persistence module $M$ of finite type if each component module is a finitely generated $\mathds{K}$-module, and if there exist positive integer $m_{0}$ such that the maps $\varphi$ are isomorphisms for $i\geq m_{0}$.
@@
From our requirements, from now on let us assume in additional that $\mathds{K}$ is a field and the above simplicial complex $K$ is finite. This means each group in chain complex 18 is a finite dimensional vector space over $\mathds{K}$. In this case, the ‘filtration’ on the chain complex 18, which will be called a persistent complex[77], is actually of finite length. In other words, the total chain complex $C(K)$ and the homology module $H(C(K),\mathds{K})$ are both persistent module of finite type. The total chain complex $C(K)$ is not only a graded $\mathds{K}$-module but also a $\mathds{K}[x]$-module with the variable $x$ of degree 1 acts as shift map via the filtration. The total homology module $H(C(K),\mathds{K})$ retains this $\mathds{K}[x]$-module structure. And from the structure theorem for principal integral domains (PID), note that $\mathds{K}[x]$ is a PID for a field $\mathds{K}$, we have the following decomposition
(21) $H(C(K),\mathds{K})\cong\mathop{\oplus}\limits_{i}x^{t_{i}}(\mathds{K}[x])\oplus(\mathop{\oplus}\limits_{j}x^{r_{j}}(\mathds{K}[x]/(x^{s_{j}}\mathds{K}[x])))$
@@
PD is one of the common visualization methods of the PH. PD is a multiset of points in the extended two-dimensional plane ¯$\overline{\mathds{R}}^{2}$ where $\overline{\mathds{R}}=\mathds{R}\cup\infty$, whose coordinates are respectively correspond to the birth and death time of the homology generators, union the diagonal points [20]. Given two PDs $P_{1}$ and $P_{2}$, there are two commonly used metrics to compare them, that is, the Wasserstein-$q$ distance $W_{q}$ and Bottleneck distance $W_{\infty}$. They are defined as following:

### v3

@@ 1.1. Primary Structures of Proteins
Protein is a class of large biomolecules, which plays important role in biochemical reactions. Amino acids (AA) are the constituent units of proteins. Specifically, each AA consists of an amino group (-$\text{NH}_{2}$), a carboxyl group (-COOH) and a side chain group (-R). These three parts are connected through a carbon atom, which we call it the $\alpha$-Carbon and denote it as $\text{C}_{\alpha}$. The type of AA is entirely determined by its side chain group. AAs form peptide bonds through dehydraton condensation, and then determine the primary structure of proteins, called the peptide chain which is shown in Fig.1. An individual AA in the peptide chain is called a residue. Each peptide chain has its backbone, which consists of amino nitrogen, $\text{C}_{\alpha}$, carboxyl carbon and carboxyl oxygen of each residue. Note that a peptide chain is naturally oriented by its formation, that is, a peptide chain always begin with a amino nitrogen and end with a carbonyl carbon.
Figure 1. A peptide chain. The backbone of this peptide chain is the set $\{\dots,\text{C}_{\alpha_{i-1}},\text{C}_{i-1},\text{O}_{i-1},\text{N}_{i},\text{C}_{\alpha_{i}},\text{C}_{i},\text{O}_{i},\dots\}.$ Blue rectangles display the peptide bonds. Here, $\text{res}_{i}$ represents the $i$-th residue of the peptide should be given in the caption.
@@ 1.2. Geometry in Biology
The geometry plays an important role in the study of the structure of proteins and even other biomolecules. Scientists found that three-dimensional atomic structures of these molecules provide rich information for understanding how these molecules carry out their biological functions. Among all these research, we found several cutting points that geometry can enter the study of the structures of proteins or biomolecules.
As commonly understood, a molecule consists of moving particles(atomic nuclei and electrons) held together by electrostatic and magnetic forces in nature, lacking definitive boundaries. However, visualizing a molecule is still a very useful chemical concept and has proved its value in medicinal chemistry.
@@
(1) $\hbox{OBSERVED OCCURRENCE}\sim\text{exp}(-\frac{\hbox{ENERGY}}{RT^{\ast}})$
@@ 1.5. Arrangement of the Paper
The reminder of this paper is organized as follows. In Section 2, we defined the truncated Hop distance to compare the structure of two discrete curves. Moreover, we explained how to treat one layer of amyloid fibril as a discrete curve. And we suggest that the effect in the formation of amyloid fibril is relatively mild based on the result of the truncated Hop distance. Preliminaries of the definitions of curvature and torsion on the discrete curves are presented in Section 3. By computing the discrete curvature and torsion on single layers of amyloid fibril, we suggest that the anomalies reflect slight affects of the layer-layer hydrogen bonds. Finally in Section 4, we introduced the distances between persistent diagrams (PDs), and showed that how these distances work in quantifying the difference of the structure of amyloid fibrils. The atomic models for crystal X-ray diffraction structures of human TTR and cryo-EM structures of ATTR wt and ATTRv were obtained from Protein Data Bank (PDB) [4].
@@ 2.2. Applications to Analysis of the Structure of Amyloid fibrils
Figure 2. Convert peptide chains into discrete curves. (A) The top is the cartoon representation of chain A of TTR V30M mutant (PDB entry: V30Mn), and the bottom is its corresponding discrete curve. (B) The top is the cartoon representation of chain A of one amyloid form of V30Mn (V30Ma-dimer), and the bottom is its corresponding discrete curve. Red balls represent the $\text{C}_{\alpha}$ atoms.
@@
Figure 3. The truncated hop distance between chain A of V30Mn and V30Ma-dimer. The entries represented by the rows and columns are the corresponding residue numbers. Here, we only considered the truncated hop distance of AAs of the same sequence fragment.
It is clear that the values near the diagonal are less than 1.5Å, which illustrates that there could not be stretching and compressing in the process of the formation of amyloid fibril. We suggest that the effect is relatively mild in the formation from chain A of V30Mn to chain A of V30Ma-dimer. Conversely, differences in hop distance greater than 25Åare primarily found in three components4A. The first and second components correspond to the hop distances between the $\beta_{E}$-strand (residues 67-73) and the $\beta_{F}$-strand (residues 88-97), and the hop distances between the EF-loop (residues 82-87) and the GH-loop (residues 113-114), respectively (Fig.4B). This is due to the unfolding of the $\alpha$-helix (residues 74-82) and the dissociation of the inter-strand and inter-loop hydrogen bond networks in well-folded TTR. The third component corresponds to the hop distances between the DE-loop (residues 56-66) and the $\beta_{H}$-strand (residues 115-123). However, this component is absent in some I84S ATTRv structures, specifically I84Sa-p3 and I84Sa-p4 (Fig.4A). The DE-loop forms a gate of the polar channel in the ATTR fibril core and varies in different polymorphs. Most ATTRwt and ATTRv fibril cores prossess a complete shape of the gate (i.e., closed gate and open gate), while the gate is incomplete or absent in some I84S mutated variants (i.e., broken gate and absent gate, Fig.4C[51].
Moreover, we identified local regions which are sensitive to the conformational transition of TTR by adopting the (truncated) $n$-hop distance. These regions contain the binding sites of Tafamidis, which is approved by the FDA in 2019 for the treatment of ATTR-caused disease. It suggests that the geometrical descriptor we defined gives insight into the drug design for stabilizers to inhibit the transition to the amyloid fibril state. More details will be shown in the biology paper [39].
Figure 4. Structural analysis based on the binary matrices with 25Åcutoff distance. (A) Binary matrices constructed by setting the cutoff distance to 25Å. (B) The sequence fragments corresponding to the first and second brighter regions and their hydrogen bond interaction. (C) The shape of gate of chain A of ATTR.
@@ 3. Discrete Curvatures and Discrete Torsions
The discrete curvature and discrete torsion based on the cross ratio and osculating circle was first introduced in [50], which is used in computer vision to estimate the differences between the background smooth curve and the discrete sample curve. The discrete curvature defined in this way is the discretization of sectional curvature on a smooth curve in a more fashion words. In the paper [39], we introduced this important parameter to the analyzing the structure of proteins. We found that compared with the other geometrical discriptors like backbone dihedral angles, the discrete curvatures and discrete torsions introduced in [10], discrete curvature resp. discrete torsion is a order two resp. order three parameter which is more sensitive to the weak interactions. Moreover it is still easy to compute although more complicated compared with the one introduced in [10] since we need to solve a $3\times 3$ linear equation. In this section, we will give a detailed introduction on these parts.
@@ 3.1. Curvatures and Torsions for Smooth Curves
(3) $0=\frac{d}{ds}\{B(s),T(s)\}=\{B^{\prime}(s),T(s)\}+\{B(s),T^{\prime}(s)\}=\{B^{\prime}(s),T(s)\}$
@@ 3.2. Basics in quaternions
Let $a,b,c,d\in\operatorname{Im}\mathds{H}$ be four points not lying on a common circle. Then, the imaginary part of the cross-ratio is the normal of the circumsphere (or plane) at $a$, i.e., for a proper circumsphere with center $m$, we have the imaginary part of $\operatorname{Im}\operatorname{cr}(a,b,c,d)$ is parallel to the vector $(m-a)$ denoted as $\operatorname{Im}\operatorname{cr}(a,b,c,d)\mathrel{/\mskip-2.5mu/}(m-a)$.
@@
We only need to prove the Equation (11) holds. From Lemma 3.4 the four points $A,B,C$ and $D$ lie on a common sphere i.e. the circumsphere of $a,b,c,d$. Let us fix an unit vector $v_{0}\in\mathds{R}^{3}$. It is clear that the space $\mathds{C}_{v_{0}}:=\{[r,kv_{0}]:r,v\in\mathds{R}\}$ is isomorphic to the complex plane $\mathds{C}$. Let $\sigma$ is the Möbius transformation that maps the circumsphere of $a,b,c,d$ to $\mathds{C}_{v_{0}}$.
@@ 3.3. Curvatures and Torsions for Discrete Curves
Let $\gamma_{i}$ be a vertex of the discrete curve. From the Proposition 3.7, for each four elements tuple $(\gamma_{i-1},\gamma_{i},\gamma_{i+1},\gamma_{i+2})$ there exist four inserting points (see Figure 5) i.e.
@@
Figure 5. The computation of discrete curvature and discrete torsion where the inserting points were marked with green and the orange circle is the osculating circle at $\gamma_{i}$.
@@ 3.4. Applications to Analysis of the Structure of Amyloid fibrils
We have shown a way of treating a peptide chain as a discrete curve and its application in the Section 2. In this section, we considered another way as following: Recall that from the Figure 1, the backbone consists of amino nitrogens, $\alpha$-carbons, carbonyl carbons and carbonyl oxygens. We treated the the coordinates of amino nitrogens, $\text{C}_{\alpha}$s, and carbonyl carbons on the backbone as the vertices of a discrete curve in $\mathds{R}^{3}$.
For a fixed layer of ATTR, we may assume its layer index in the protein is $\ast$, let $\text{N}^{\ast}_{i}$ (resp. $\text{C}^{\ast}_{\alpha_{i}}$ and $\text{C}^{\ast}_{i}$) be the coordinates of amino nitrogens (resp. alpha carbons and carbonyl carbons) with $i\in\mathds{Z}^{\geq 1}$ being the index of amino acids (AAs) in the layer. Therefore, the curvature and torsion at each atom could be derived from the position of itself, one atom ahead and two atom after, e.g. $\text{C}^{\ast}_{\alpha_{i}}$, $\text{C}^{\ast}_{i}$, $\text{N}^{\ast}_{i+1}$ and $\text{C}^{\ast}_{\alpha_{i+1}}$ determine the curvature and the torsion at $\text{C}^{\ast}_{i}$. We computed the absolute values of the curvatures ($|\kappa|$) and torsions ($|\tau|$) at different atoms from all residues (the curvatures resp. torsions and atoms involved in the computation are listed in 17) in each ATTR fibril structure, as well as the arithmetic mean of absolute values of the curvatures ($\overline{|\kappa|}$) and torsions ($\overline{|\tau|}$), and the (unbiased) variance of $|\kappa|$ ($S^{2}(|\kappa|)$) and $|\tau|$ ($S^{2}(|\tau|)$), respectively.
(17) $\text{C}^{\ast}_{i-1},\text{N}^{\ast}_{i},\text{C}^{\ast}_{\alpha_{i}},\text{C}^{\ast}_{i}$$\to$ $\hbox{Curvatures and Torsions at }\text{N}_{i}$
$\text{N}^{\ast}_{i},\text{C}^{\ast}_{\alpha_{i}},\text{C}^{\ast}_{i},\text{N}^{\ast}_{i+1}$$\to$ $\hbox{Curvatures and Torsions at }\text{C}_{\alpha_{i}}$
$\text{C}^{\ast}_{\alpha_{i}},\text{C}^{\ast}_{i},\text{C}^{\ast}_{i+1},\text{C}^{\ast}_{i+1}$$\to$ $\hbox{Curvatures and Torsions at }\text{C}_{i}$
The pseudo-code of computing curvature and torsion is shown as the Algorithm 1.To solve the curvature and torsion at one atom, one needs to solve a $3\times 3$ linear equation for determining the center of the osculating circle. The computation complexity is a constant. Therefore the computation complexity for computing all discrete curvatures and torsions is linear with respect to the number of atoms. Moreover the computation could be conduct parallelly for large size of data set.
@@
Let (1)($\ast$-1)( ∗ - 1 ) and (+1)($\ast$+1)( ∗ + 1 ) be two nearby chains of the $\ast$-th chain. The backbone hydrogen bonds were always between accepter, the carbonyl oxygen of the $i$-th AA of the chain $\ast$ (denoted by $\text{O}_{i}^{\ast}$), and doner, the amino nitrogen of $(i+1)$-th AA of the chain (1)($\ast$-1)( ∗ - 1 ) or the chain (+1)($\ast$+1)( ∗ + 1 ) (denoted by $\text{N}_{i+1}^{(\ast-1)}$ or $\text{N}_{i+1}^{(\ast+1)}$, respectively), according to the direction of the carbonyl. Let $\tilde{d}_{i,+}^{\ast}$ (resp. $\tilde{d}_{i,-}^{\ast}$) be the distance between $\text{O}_{i}^{\ast}$ and $\text{N}_{i+1}^{(\ast+1)}$ (resp. the distance between $\text{O}_{i}^{\ast}$ and $\text{N}_{i+1}^{(\ast-1)}$). The squared distance differences at the $i$-th AA of the chain $\ast$ is defined as,
(18) $\tilde{d}_{i}^{\ast}:=|(\tilde{d}_{i,-}^{\ast})^{2}-(\tilde{d}_{i,+}^{\ast})^{2}|$
(See Fig.6).
Figure 6. The squared distance differences at the $i$-th AA of the chain B. The atoms are marked in different colors, i.e., carbons marked with grey, nitrogens marked with blue and oxygens marked with red.
Table 2. The mean and variance of absolute values of curvatures and torsions for V30Ma-dimer.
$\bar{\kappa}$ $\bar{\kappa}_{N}$ $\bar{\kappa}_{C\alpha}$ $\bar{\kappa}_{C}$ $\bar{\tau}$ $\bar{\tau}_{N}$ $\bar{\tau}_{C\alpha}$ $\bar{\tau}_{C}$ $S^{2}(\kappa)$ $S^{2}(\kappa_{N})$ $S^{2}(\kappa_{C\alpha})$ $S^{2}(\kappa_{C})$ $S^{2}(\tau)$ $S^{2}(\tau_{N})$ $S^{2}(\tau_{C\alpha})$ $S^{2}(\tau_{C})$
V30Ma-dimer 0.305 0.445 0.421 0.050 5.048 1.521 1.858 11.766 0.052 0.024 0.034 0.000 64.589 1.441 9.727 114.901
We conducted the computation to several amyloid fibrils. We found that the torsions at carbonyl carbons are significantly larger than the one in other places (we list one example V30Ma-dimer in Figure 7 (A) and Table 2). We showed that this anomalies reflect slight affects of the layer-layer hydrogen bonds.
Note that the carbonyls will form hydrogen bonds with amino hydrogen in the nearby chain along fibril axis, which play important roles in the stabilization of amyloid fibrils. We suggested that this anomaly in torsion reflects the existence and power of hydrogen bonds. We made the correlation analysis for each ATTR by using the linear model between the absolute value of torsions at carbonyl carbons in the $i$-th AA at the $\ast$-th layer ($\tau_{i}^{\ast}$) and the squared distance differences ($\tilde{d}_{i}^{\ast}$), where $\ast$ runs over all possible layers and $i$ run over all AAs in the layer (the case V30M-dimer is showed in Fig.7 (B)). The zero assumption is listed as following,
(19) $H_{0}:\hbox{The absolute value of torsion }|\tau_{i}^{\ast}|\hbox{has non-negative correlation with }\tilde{d}_{i}^{\ast}$
Note that the $p$-value is computed using the Wald Test with the $t$-distribution of the test statistic and S.D stands for the standard error.
We found the $p$-value of $H_{0}$ is rather small (We show one example V30Ma-dimer in Table 3). Therefore, we suggested that $|\tau_{i}^{\ast}|$ is negatively correlated with $\tilde{d}_{i}^{\ast}$. Since the strength of hydrogen bonds is negatively related with the square of the distance, we could derive from all these facts that the above anomaly of the torsion is caused by the layer-to-layer hydrogen bonds.
[…]
Figure 7. (A) Distributions of absolute values of Torsions for V30Ma-dimer. The $x$-axis represents the absolute value of torsion and the $y$-axis represents the number of atoms. We take purple bar for nitrogens, orange for alpha carbons and green for carbonyl carbons. The discrete torsion is dimensionless. (B)The scatter of absolute values of torsions at carbonyl carbons $\tau_{i}^{\ast}$ against the squared distance differences $\tilde{d}_{i}^{\ast}$ between nearby layers for all amyloid fibrils of V30Ma-dimer. Recall that $\tilde{d}_{i}^{\ast}$ is defined in Equation 18 where $\ast$ stands for the index of the layer in the protein and $i$ stands for the index of AA in the layer. The horizontal axis stands for the value $\tilde{d}_{i}^{\ast}$, the vertical axis stands for $|\tau_{i}^{\ast}|$ and the dashed grey line is the result of regression.
Table 3. The correlation of $|\tau_{i}|$ and $|(\tilde{d}^{i}_{\cdot,-})^{2}-(\tilde{d}^{i}_{\cdot,+})^{2}|$ for V30Ma-dimer. S.D stands for the standard error
@@ 4.1. Persistent Homology in a Nutshell
where $C_{n}(K)$ is the free $\mathds{K}$-module generated by the basis of simplicies $\sigma=[x_{0},\cdots,x_{n}]$ in $K$ of length $n+1$ and the boundary map $\partial_{n+1}:C_{n+1}(K)\to C_{n}(K)$ is a $\mathds{K}$-linear map defined in a natural way.
@@
where $Z_{n}^{t}$ is the $n$-th cycle group of $K_{t}$, i.e., $\ker\partial_{n}^{t}$, and $B_{n}^{t^{\prime}}$ is the $n$-th boundary group of $K_{t^{\prime}}$, i.e., $\operatorname{Im}\partial_{n+1}^{t^{\prime}}$. The rank of $H_{n}^{t\to t^{\prime}}$ is called the $(t,t^{\prime})$-persistent $n$-th Betti number denoted as $\beta_{n}^{t\to t^{\prime}}$ [80].
@@
A persistence $\mathds{K}$ module $M$ is a family of $\mathds{K}$-modules $M_{i}$, together with homomorphisms $\varphi_{i}:M_{i}\to M_{i+1}$. We say a persistence module $M$ of finite type if each component module is a finitely generated $\mathds{K}$-module, and if there exist positive integer $m_{0}$ such that the maps $\varphi_{i}$ are isomorphisms for $i\geq m_{0}$ [80].
@@
From our requirements, from now on let us assume in additional that $\mathds{K}$ is a field and the above simplicial complex $K$ is finite. This means each group in chain complex 20 is a finite dimensional vector space over $\mathds{K}$. In this case, the ‘filtration’ on the chain complex 20, which will be called a persistent complex [80], is actually of finite length. In other words, the total chain complex $C_{\bullet}(K)$ and the homology module $H_{\bullet}(C_{\bullet}(K),\mathds{K})$ are both persistent module of finite type. The total chain complex $C_{\bullet}(K)$ is not only a graded $\mathds{K}$-module but also a $\mathds{K}[x]$-module with the variable $x$ of degree 1 acts as shift map via the filtration. The total homology module $H_{\bullet}(C_{\bullet}(K),\mathds{K})$ retains this $\mathds{K}[x]$-module structure. And from the structure theorem for principal ideal domains (PID), note that $\mathds{K}[x]$ is a PID for a field $\mathds{K}$, we have the following decomposition
(23) $H_{\bullet}(C_{\bullet}(K),\mathds{K})\cong\mathop{\oplus}\limits_{i}x^{t_{i}}(\mathds{K}[x])\oplus(\mathop{\oplus}\limits_{j}x^{r_{j}}(\mathds{K}[x]/(x^{s_{j}}\mathds{K}[x])))$
@@
PD is one of the common visualization methods of the PH. PD is a multiset of points in the extended two-dimensional plane ¯$\overline{\mathds{R}}^{2}$ where $\overline{\mathds{R}}=\mathds{R}\cup\{\infty\}$, whose coordinates are respectively correspond to the birth and death time of the homology generators, union the diagonal points [22]. Given two PDs $P_{1}$ and $P_{2}$, there are two commonly used metrics to compare them, that is, the Wasserstein-$q$ distance $W_{q}$ and Bottleneck distance $W_{\infty}$. They are defined as following: